OpenStudy (anonymous):

Medals for you! can i have some help from the people at open study? solve each equation 1.2cos^2(theta)-3cos(theta) +1=0 , 0

OpenStudy (anonymous):

@SithsAndGiggles would you be able to help me?

OpenStudy (anonymous):

It's a quadratic in disguise. Try writing \(x=\cos\theta\) to make it easier to see: \[2x^2-3x+1=0\]

OpenStudy (anonymous):

wow thats so much clearer...good!

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

Not quite, you can factor the expression: \[2x^2-3x+1=(2x-1)(x-1)=0\] which gives \(x=1\) *and* \(x=\dfrac{1}{2}\). Remember that \(x=\cos\theta\), so in fact you have \[\cos\theta=1~~\text{and}~~\cos\theta=\frac{1}{2}\]

OpenStudy (anonymous):

right, which makes a lot more sense..this is great, honestly thanks for helping me with such a simple one! thanks XD